OpenStudy (anonymous):

If i have -4.5x - (-1.5) + 6.5x = -12.5 and i want to combine the x's, should i do -4.5x - 6.5x or -4.5x + 6.5x?

OpenStudy (anonymous):

the second thing you wrote... -4.5x + 6.5x

OpenStudy (anonymous):

you can rearrange the expression however you want, just keep track of the + or - in front of each term. How would you combine the -(-1.5) and the -12.5 on the opposite side of the equation?

OpenStudy (anonymous):

Ok. Thanks! Um... I dont know. Probably subtract them. I havent gotten that far into the equation yet. But thats a good question.

OpenStudy (anonymous):

Post what you get... I will check it.

OpenStudy (anonymous):

Ok. I have now, 2x - (-1.5) = -12.5

OpenStudy (anonymous):

good so far

OpenStudy (anonymous):

Ok so what do i do to get 2x alone on one side?

OpenStudy (anonymous):

Wait. WOuld i get 2x = -14?

OpenStudy (anonymous):

maybe this will help... if I said I have 2x - STUFF = -12.5, that means I am subtracting "STUFF" from the left side. What if I add "STUFF" to both sides? On the left side, 2x - STUFF + STUFF <<-- note: -STUFF + STUFF = ZERO!!! and on the right, -12.5 + STUFF So it would be 2x = -12.5 + STUFF In your problem, that STUFF is (-1.5)

OpenStudy (anonymous):

Now i have x = -7

OpenStudy (anonymous):

yeah, you add (-1.5) to both sides, so on the right, it's -14. And 2x = -14 --->>> x = -7

OpenStudy (anonymous):

Cool! Thanks!!

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

Now i have y = -0.75 + 3.25x. When i solved it i got y = -23.5. Is that correct?

OpenStudy (anonymous):

Uh, you can't solve it without being given an x value... did it say solve for y given some specific x?

OpenStudy (anonymous):

Oh sorry. x = -7

OpenStudy (anonymous):

that's what I get too :) Looks good :)

OpenStudy (anonymous):

Ok but when i plug it all in, it doesnt make sense. Im doing a solve by substitution problem. If you want, i can post it.

OpenStudy (anonymous):

sure, I was just helping with the parts, but seeing it all would help